{JC, Paco, Yehonathan, Ian}
POY lens was moved to infront of the POY steering mirror to make the POU beam focused on the POY11 RFPD. We measured the DC output with an oscilloscope and optimized it with the steering mirrors. We get ~ 16.5mV.
The new lens position blocked the BS OpLev ingoing beam, so we repositioned the OpLev mirrors to make the beam path not hit the lens.
We went to the control room to observe the PDH signal. We observed a series of PDF osscillation and then the signal died infront of our eyes! There is just noise.
We go and check the +/-15V powering the RFPD and we find that the V- is ~ 14V which is good but the V+ was ~ 2.7V which is not.
We went to the PD interface and measure the POY11 output oltages using a breakout board and got the same result.
The PD interface was taken out for inspection. All the OP27 on channel 3 were replaced with new ICs (without need turns out)...
The PD interface card turned out to be OK. What happened is that one of the Kepcos in the RF rack died because its fan crumbled as seen in Attachment #2 (could this be the source of burning smell?). In response, the rack was drawing from the other Kepco (connected in parallel) way too much current (4A) and the current limiter dropped its voltage from 15V to 2.7V.
The Kepco pair was removed and replaced with a single Sorensen. The POY PDH signal was restored (see attachment). |